OpenStudy (anonymous):

What is the solution to the system of equations x+y=2 x-y=4

OpenStudy (anonymous):

Add both equation. you will get value of x the substitute it in any eq. for y

OpenStudy (anonymous):

Add both equations (x+y)+(x-y) = 2 + 4 x + y + x - y = 6 2x = 6 x = 3 Now substitute x=3 in either of the given equations In x-y = 4 3 - y =4 y =3-4 y =-1

OpenStudy (anonymous):

yea das tight x=3 y=-1

OpenStudy (anonymous):

thank you